OpenStudy (anonymous):

simplify 6 fourth root of 4 minus 7 cubed root of 4 minus3 fourth root of 4

OpenStudy (zepp):

\[\large \frac{6}{4}\sqrt{4}-7\sqrt[3]{4}-\frac{3}{4}\sqrt[4]{4}\]

OpenStudy (anonymous):